The product of a positive even integer and the reciprocal of the next larger positive even integer equals the reciprocal of the first integer. Find the integer.
Answer by mananth(16946)

The product of a positive even integer and the reciprocal of the next larger positive even integer equals the reciprocal of the first integer. Find the integer.
Let one number be x
the other number will be x+2
reciprocal of this number = 1/ x+2
x* 1/ x+2 = 1/x
x/ x+2 =1/x
x^2=x+2
x^2-x-2=0
x^2-2x+x-2=0
x(x-2)+1(x-2)=0
(x+1)(x-2)=0
x=-1 or x=2
Answer 2